Need help with:
7/2x^2+3=5
That is 7 over 2x squared if there is any confusion.

Answer by Alan3354(69443)

7/2x^2+3=5
7/2x^2 = 2
2x^2 = 7/2
x^2 = 7/4
x = ± sqrt(7)/2